# Quantum-Inspired Ising Machines for Quantum Chemistry Calculations

### v2

@@ 1 Introduction
Quantum computing offers a transformative alternative, as originally envisioned by Feynman 3. Quantum Phase Estimation (QPE) promises FCI-level accuracy 4 but requires fault-tolerant quantum computers (FTQC). Advances by Jones et al. 5 established protocols (Programmable Ancilla Rotations (PARs), Phase Kickback, Fault-Tolerant Gate Approximation Sequences, etc.) for scalable ab initio simulations, while hybrid quantum-classical algorithms such as the Variational Quantum Eigensolver (VQE) 6, 7 emerged to address Noisy Intermediate-Scale Quantum (NISQ) era constraints. VQE reduces circuit depth by combining parameterized quantum circuits with classical optimization, but faces challenges including ansatz design 8, barren plateaus 9, and measurement overhead 10.
Numerous improvements, such as UCCSD 11, 12, 13, 14, 15, 16, 17, ADAPT-VQE 18, k-UpCCGSD 19, Qubit-ADAPT-VQE 20, and Quantum Subspace Expansion (QSE) 22, have been developed, along with hardware-efficient architectures and quantum machine learning extensions 23. Real hardware implementation of these algorithms based on photonics 6, 24, superconducting qubits 25, 26, 28, 27, and trapped ions 29, 30, highlighting both promise and limitations. Beyond gate-based approaches, real-space, real-time simulations offer direct dynamical insights at high computational cost 31, 32, 21, 33, 34, 35.
Adiabatic quantum computing and quantum annealing also provide alternative formulations. Although limited by hardware connectivity, the mapping of the electronic structure to the Ising or Quadratic unconstrained binary optimization (QUBO) Hamiltonians has been demonstrated on D-Wave devices 36, 37, 38. Hybrid strategies, such as Qubit Coupled Cluster (QCC) 39 and Quantum Annealer Eigensolver (QAE) 40, extend these approaches to intermediate-size molecular systems. Collectively, these efforts underscore both the potential and current limitations of quantum algorithms for chemistry.
Recently, Quantum Annealing-Inspired Algorithms (QAIAs) have emerged as scalable, high-performance alternatives. While not inherently quantum, they leverage principles of quantum annealing to achieve more than $98\%$ accuracy in Ising ground-state estimation. Two leading approaches are Coherent Ising Machines (CIMs), based on networks of quantum optical parametric oscillators 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, and Simulated Bifurcation (SB) algorithms, which exploit nonlinear bifurcation dynamics of Kerr-nonlinearity for efficient optimization 51, 52, 53. CIMs benefit from all-to-all optical coupling and large-scale operation at room temperature, while SB algorithms offer computationally efficient variants successfully applied to molecular Hamiltonians.
@@ 2.1 Mapping the Molecular Hamiltonian into Ising Hamiltonian
Mapping a molecular Hamiltonian into an Ising-type Hamiltonian is a multistep procedure that connects ab initio quantum chemistry with quantum optimization and quantum annealing methods. A detailed description of this framework is provided by Xia et al. 37. The process begins with the construction of molecular orbitals, typically from a chosen atomic orbital basis, such as STO6GSTO-6G or larger basis sets, depending on the required accuracy. Spin orbitals are then obtained by combining spatial orbitals with spin functions while also accounting for the overlap integrals between atomic orbitals.
@@ 2.2 Ising Machines
Several approaches have been proposed for sampling data to feed into Ising solvers for molecular electronic structure calculations. Previous studies have employed D-Wave annealers for this purpose 36, 37, 38. In this work, we employ quantum-inspired Ising solvers, specifically the CIM and the SB algorithm, which offer distinct advantages over D-Wave–based quantum annealing approaches. Both CIMs and SBs have demonstrated high efficiency in determining the minimum energy configurations of the Ising Hamiltonian, often achieving superior success probabilities. The operational principles and algorithmic structures of these solvers are described in detail in the following section.
From a physical standpoint, the operation of CIMs can be understood in terms of oscillator dynamics: each oscillator undergoes a parametric gain that amplifies the field once the pump exceeds the threshold, while nonlinear loss (saturation) constrains the amplitude to one of two stable phase states 74. The couplings between oscillators are implemented by modulating the pump or injection signals according to the Ising coupling matrix $J_{ij}$, thereby encoding the global Ising energy into the gain–loss landscape of the entire network. Quantum and classical noise seed the initial phases, and the ensuing nonlinear dynamics guide the system toward low-energy phase configurations. Thus, CIMs exploit the physics of parametric oscillation and nonlinear bifurcation to identify the ground-state configuration.
Previous studies have proposed a range of quantum-inspired algorithms based on CIMs that address the Ising problem through the Heisenberg–Langevin formalism, most notably the Chaotic Amplitude Control (CAC), Chaotic Feedback Control (CFC), and Separated Feedback Control (SFC) algorithms 48, 75.
@@
Beyond CIM variants, the discrete Simulated Bifurcation (dSB) algorithm provides a highly efficient quantum-inspired framework for solving Ising problems. Rooted in the bifurcation theorem and the adiabatic evolution of nonlinear Hamiltonian systems 88, the dSB algorithm is particularly noted for its parallel processing capabilities and its ability to rapidly locate local minima near the ground state 51, 52. Unlike the standard ballistic formulation, the dSB variant enhances optimization performance by discretizing the variables within the interaction term. The time evolution of the spins is governed by:
@@
To further improve the solutions obtained from the Ising solvers, namely, the CIMs and SB, we apply a post-processing technique based on a greedy algorithm, specifically the steepest descent method provided by the SteepestDescentSolver in D-Wave’s Ocean software 89. This method iteratively refines the spin configurations returned by the quantum processing unit (QPU), aiming to minimize the energy of the system and identify lower-energy configurations. In essence, once a local ground state is obtained from the Ising solver, the algorithm explores its neighboring configurations to ensure that the best possible local minimum is achieved.
@@
Figure 2: Schematic of the hybrid algorithm employed in this study, extending and refining the framework originally introduced by Alán Aspuru-Guzik 10.
@@ 3 Results
To assess the accuracy of our algorithms, we compare their results with those obtained from standard electronic structure methods, namely the Hartree-Fock (HF) and Complete Active Space Configuration Interaction (CASCI) methods. The HF method represents the molecular wavefunction as a single Slater determinant and provides the variationally optimal solution within the mean-field approximation, whereas the CASCI method yields the exact solution for the chosen basis set and active space and becomes equivalent to the full configuration interaction (FCI) method when all spin-orbitals are included in the active space 76, 77. These reference methods, therefore, serve as benchmarks to evaluate the performance and accuracy of the proposed algorithms.
@@ 3.1 Single-shot
Fig. 3 illustrates the calculated ground-state energy of the hydrogen molecule as a function of its internuclear bond length. The results are presented for four quantum-inspired algorithms: SFC, CFC, dSB, and CAC. Each subfigure displays the corresponding energy profile, where the different colored lines represent variations in $r$, ranging from $r=2$ to $r=6$, a variational parameter introduced in the original paper 37.
@@
Across all four algorithms, a consistent trend characteristic of molecular dissociation curves is observed. As the internuclear bond length increases from very short distances, the energy initially decreases, reaches a minimum, and then gradually increases, eventually approaching a plateau at larger bond lengths. This minimum energy corresponds to the equilibrium bond length and the ground-state energy of the molecule. The parameter $r$ influences both the accuracy and convergence of the energy estimation, with certain values of $r$ exhibiting closer agreement with the expected energy profile, particularly near the equilibrium bond length. In the XBK method, the energy plotted as a function of $r$ represents the collective ground-state energy of the molecule, as the transformation encodes the full molecular Hamiltonian into an expanded qubit space, capturing all interatomic interactions simultaneously rather than individual bond energies 37, 77.
@@ 3.2 Multi-shot
Table 2 summarizes the measured and estimated per-evaluation TTS for small-molecule ground-state energy calculations across gate-model VQE, annealer-based XBK/QCC, and quantum-inspired Ising-machine simulations. Two gate-model experiments on real hardware illustrate the wide range in end-to-end runtime: a superconducting IBMQ Manila run required approximately $334\pm 40$ s per energy evaluation, while an ion-trap experiment on AQT Marmot reported much longer durations of $13297\pm 625$ s, both including queueing, compilation, execution, and readout overheads 78.
Annealer-based approaches (XBK/QCC) executed on D-Wave yield faster single-evaluation times in the 1-9 seconds cite Teplukhin2021; however, they include several computationally expensive components, such as quadratization, embedding, and multiple qbsolv 80 subcalls (qbsolv is a Python library utilized as a decomposition solver), and both classical and quantum contributions scale rapidly with system size.
By contrast, the quantum-inspired Ising-machine simulations proposed in this work, achieve short wall-clock times per energy evaluation, 131-3 s on GPUs. In particular, total potential energy computation for H2 required only $1.2$ s per bond-length parameter $r$ (Fig. 4), while for H2O the simulator required $2.5$ s per parameter $r$ (Fig. 5). Crucially, this highlights a distinct practical advantage of the GPU-based Ising approach by bypassing the queueing latencies and programming overheads inherent to current cloud-based QPUs; we achieve superior computational throughput and immediate scalability.
@@
H2 VQE 78 IBMQ Manila $334\pm 40$ s End-to-end TTS includes queue, compilation, execution, and readout.
H2 VQE 78 AQT Marmot $13{,}297\pm 625$ s End-to-end TTS includes queue, compilation, execution, and readout.
H2O XBK (total) 79 D-Wave more than 8 s (est.) Estimated from JCP Fig. 3 (time vs. pre-quadratization qubits).
H2O QCC (total) 79 D-Wave more than 6 s (est.) Estimated from Fig. 3.
@@
H2O Quantum-Inspired Ising Machine GPU 2.5 s Proposed efficient simulation using the Ising Machine approach.
@@ 4 Conclusion
Our numerical analysis, consistent with recent studies that compare the performance of QAIA 81, indicates that CFC variant exhibits superior efficiency compared to other CIM-inspired algorithms 81, 82.
@@
It is noteworthy that our proposed GPU-based algorithm for quantum chemistry calculations can efficiently generate large numbers of samples, enabling the treatment of large-scale scientific problems with minimal time and resource overhead due to its inherent parallelism 48, 47, 51, 52. This high-throughput sampling not only facilitates exploration of Ising-form problems but also provides a powerful framework for investigating electron configurations in molecular systems to identify their stable states. With this physics-inspired, GPU-accelerated approach, the exploration of electronic structures in complex molecules can be performed more efficiently, offering promising applications in material design—such as optimizing catalytic surfaces, studying superconducting materials, and engineering novel semiconductors—and in drug discovery, for instance, in predicting binding affinities, screening molecular conformations, and modeling reaction pathways in biomolecular systems.
@@
Most importantly, in contrast to purely digital optimization approaches, the physical realization of CIMs offers an analog computational paradigm with distinct advantages for quantum chemistry applications. CIMs evolve continuous dynamical variables, such as optical field amplitudes, thereby avoiding the numerical precision limitations and significant memory overhead associated with high-accuracy digital computation. Furthermore, intrinsic quantum noise in optical parametric oscillator networks provides a physically meaningful source of stochasticity, enabling efficient exploration of complex energy landscapes and reducing the likelihood of stagnation in local minima 83. Although the present study employed SimCIM, it is noteworthy that hardware-realized CIMs are expected to offer additional advantages, including faster physical convergence, enhanced scalability, and improved sampling fidelity due to genuine quantum fluctuations. When combined with quantum-inspired algorithms such as CFC, SFC, CAC, and dSB variants as demonstrated in this work, CIM-based hybrid solvers present a scalable and physics-informed alternative to both classical heuristics and hardware-restricted quantum annealers. Their demonstrated ability to approximate molecular ground-state energies using modest computational resources highlights their potential as a key component of future analog–digital hybrid computing frameworks in quantum chemistry.
@@
During the preparation of this work, Takesue et al. 84 demonstrated that a CIM based on degenerate optical parametric oscillators can efficiently solve classically hard NP problems, such as finding maximum independent sets in large graphs, by mapping them onto an Ising Hamiltonian and exploiting the machine’s intrinsic parallelism and energy minimization dynamics. This success story provides a testimony that photonic-based, quantum-inspired systems can efficiently explore complex energy landscapes, a capability directly relevant to quantum chemistry, where molecular Hamiltonians can be encoded into Ising-type models to determine ground-state energies. Leveraging CIMs thus enables systematic exploration of electronic configurations, outperforming traditional digital algorithms in variational molecular simulations.
@@
Building on this foundation, future research can enhance convergence and accuracy through optimized parameter tuning, noise mitigation, and adaptive sampling, while integration with machine learning may enable automated optimization. Embedding these methods in hybrid quantum–classical workflows could further improve efficiency. Extending the approach to larger, chemically relevant systems, including complex biomolecules, remains an important opportunity, and the hybrid algorithm could also be adapted to compute excited-state energy profiles 85, broadening its applicability in quantum chemistry, materials science, and drug discovery.

### v3

@@ 1 Introduction
Quantum computing offers a transformative alternative, as originally envisioned by Feynman 3. Quantum Phase Estimation (QPE) promises FCI-level accuracy 4 but requires fault-tolerant quantum computers (FTQC). Advances by Jones et al. 5 established protocols (Programmable Ancilla Rotations (PARs), Phase Kickback, Fault-Tolerant Gate Approximation Sequences, etc.) for scalable ab initio simulations, while hybrid quantum-classical algorithms such as the Variational Quantum Eigensolver (VQE) 6, 7, 21 emerged to address Noisy Intermediate-Scale Quantum (NISQ) era constraints. VQE reduces circuit depth by combining parameterized quantum circuits with classical optimization, but faces challenges including ansatz design 8, barren plateaus 9, and measurement overhead 10.
Numerous improvements, such as UCCSD 11, 12, 13, 14, 15, 16, 17, ADAPT-VQE 18, k-UpCCGSD 19, Qubit-ADAPT-VQE 20, and Quantum Subspace Expansion (QSE) 23, have been developed, along with hardware-efficient architectures and quantum machine learning extensions 24. Real hardware implementation of these algorithms based on photonics 6, 25, superconducting qubits 26, 27, 29, 28, and trapped ions 30, 31, highlighting both promise and limitations. Beyond gate-based approaches, real-space, real-time simulations offer direct dynamical insights at high computational cost 32, 33, 22, 34, 35, 36.
Adiabatic quantum computing and quantum annealing also provide alternative formulations. Although limited by hardware connectivity, the mapping of the electronic structure to the Ising or Quadratic unconstrained binary optimization (QUBO) Hamiltonians has been demonstrated on D-Wave devices 37, 38, 39. Hybrid strategies, such as Qubit Coupled Cluster (QCC) 40 and Quantum Annealer Eigensolver (QAE) 41, extend these approaches to intermediate-size molecular systems. Collectively, these efforts underscore both the potential and current limitations of quantum algorithms for chemistry.
Recently, Quantum Annealing-Inspired Algorithms (QAIAs) have emerged as scalable, high-performance alternatives. While not inherently quantum, they leverage principles of quantum annealing to achieve more than $98\%$ accuracy in Ising ground-state estimation. Two leading approaches are Coherent Ising Machines (CIMs), based on networks of quantum optical parametric oscillators 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, and Simulated Bifurcation (SB) algorithms, which exploit nonlinear bifurcation dynamics of Kerr-nonlinearity for efficient optimization 52, 53, 54. CIMs benefit from all-to-all optical coupling and large-scale operation at room temperature, while SB algorithms offer computationally efficient variants successfully applied to molecular Hamiltonians.
@@ 2.1 Mapping the Molecular Hamiltonian into Ising Hamiltonian
Mapping a molecular Hamiltonian into an Ising-type Hamiltonian is a multistep procedure that connects ab initio quantum chemistry with quantum optimization and quantum annealing methods. A detailed description of this framework is provided by Xia et al. 38. The process begins with the construction of molecular orbitals, typically from a chosen atomic orbital basis, such as STO6GSTO-6G or larger basis sets, depending on the required accuracy. Spin orbitals are then obtained by combining spatial orbitals with spin functions while also accounting for the overlap integrals between atomic orbitals.
@@ 2.2 Ising Machines
Several approaches have been proposed for sampling data to feed into Ising solvers for molecular electronic structure calculations. Previous studies have employed D-Wave annealers for this purpose 37, 38, 39. In this work, we employ quantum-inspired Ising solvers, specifically the CIM and the SB algorithm, which offer distinct advantages over D-Wave–based quantum annealing approaches. Both CIMs and SBs have demonstrated high efficiency in determining the minimum energy configurations of the Ising Hamiltonian, often achieving superior success probabilities. The operational principles and algorithmic structures of these solvers are described in detail in the following section.
From a physical standpoint, the operation of CIMs can be understood in terms of oscillator dynamics: each oscillator undergoes a parametric gain that amplifies the field once the pump exceeds the threshold, while nonlinear loss (saturation) constrains the amplitude to one of two stable phase states 75. The couplings between oscillators are implemented by modulating the pump or injection signals according to the Ising coupling matrix $J_{ij}$, thereby encoding the global Ising energy into the gain–loss landscape of the entire network. Quantum and classical noise seed the initial phases, and the ensuing nonlinear dynamics guide the system toward low-energy phase configurations. Thus, CIMs exploit the physics of parametric oscillation and nonlinear bifurcation to identify the ground-state configuration.
Previous studies have proposed a range of quantum-inspired algorithms based on CIMs that address the Ising problem through the Heisenberg–Langevin formalism, most notably the Chaotic Amplitude Control (CAC), Chaotic Feedback Control (CFC), and Separated Feedback Control (SFC) algorithms 49, 76.
@@
Beyond CIM variants, the discrete Simulated Bifurcation (dSB) algorithm provides a highly efficient quantum-inspired framework for solving Ising problems. Rooted in the bifurcation theorem and the adiabatic evolution of nonlinear Hamiltonian systems 89, the dSB algorithm is particularly noted for its parallel processing capabilities and its ability to rapidly locate local minima near the ground state 52, 53. Unlike the standard ballistic formulation, the dSB variant enhances optimization performance by discretizing the variables within the interaction term. The time evolution of the spins is governed by:
@@
To further improve the solutions obtained from the Ising solvers, namely, the CIMs and SB, we apply a post-processing technique based on a greedy algorithm, specifically the steepest descent method provided by the SteepestDescentSolver in D-Wave’s Ocean software 90. This method iteratively refines the spin configurations returned by the quantum processing unit (QPU), aiming to minimize the energy of the system and identify lower-energy configurations. In essence, once a local ground state is obtained from the Ising solver, the algorithm explores its neighboring configurations to ensure that the best possible local minimum is achieved.
@@
Figure 2: Schematic of the hybrid algorithm employed in this study, extending and refining the framework originally introduced by Yudong Cao et. al. 10.
@@ 3 Results
To assess the accuracy of our algorithms, we compare their results with those obtained from standard electronic structure methods, namely the Hartree-Fock (HF) and Complete Active Space Configuration Interaction (CASCI) methods. The HF method represents the molecular wavefunction as a single Slater determinant and provides the variationally optimal solution within the mean-field approximation, whereas the CASCI method yields the exact solution for the chosen basis set and active space and becomes equivalent to the full configuration interaction (FCI) method when all spin-orbitals are included in the active space 77, 78. These reference methods, therefore, serve as benchmarks to evaluate the performance and accuracy of the proposed algorithms.
@@ 3.1 Single-shot
Fig. 3 illustrates the calculated ground-state energy of the hydrogen molecule as a function of its internuclear bond length. The results are presented for four quantum-inspired algorithms: SFC, CFC, dSB, and CAC. Each subfigure displays the corresponding energy profile, where the different colored lines represent variations in $r$, ranging from $r=2$ to $r=6$, a variational parameter introduced in the original paper 38.
@@
Across all four algorithms, a consistent trend characteristic of molecular dissociation curves is observed. As the internuclear bond length increases from very short distances, the energy initially decreases, reaches a minimum, and then gradually increases, eventually approaching a plateau at larger bond lengths. This minimum energy corresponds to the equilibrium bond length and the ground-state energy of the molecule. The parameter $r$ influences both the accuracy and convergence of the energy estimation, with certain values of $r$ exhibiting closer agreement with the expected energy profile, particularly near the equilibrium bond length. In the XBK method, the energy plotted as a function of $r$ represents the collective ground-state energy of the molecule, as the transformation encodes the full molecular Hamiltonian into an expanded qubit space, capturing all interatomic interactions simultaneously rather than individual bond energies 38, 78.
@@ 3.2 Multi-shot
Table 2 summarizes the measured and estimated per-evaluation TTS for small-molecule ground-state energy calculations across gate-model VQE, annealer-based XBK/QCC, and quantum-inspired Ising-machine simulations. Two gate-model experiments on real hardware illustrate the wide range in end-to-end runtime: a superconducting IBMQ Manila run required approximately $334\pm 40$ s per energy evaluation, while an ion-trap experiment on AQT Marmot reported much longer durations of $13297\pm 625$ s, both including queueing, compilation, execution, and readout overheads 79. In addition, a quite recent work on superconductor quantum computer for finding the ground state of molecule reported that the time to find a precise ground state took about 1.5 hours.
Annealer-based approaches (XBK/QCC) executed on D-Wave yield faster single-evaluation times in the 1-9 seconds 78, 80. However, they include several computationally expensive components, such as quadratization, embedding, and multiple qbsolv 81 subcalls (qbsolv is a Python library utilized as a decomposition solver), and both classical and quantum contributions scale rapidly with system size. The estimated runtime corresponds to the end-to-end wall-clock time on the D-Wave quantum annealer, including fixed programming overhead and per-shot readout overhead. The actual quantum annealing time per run is on the order of microseconds and contributes only a negligible fraction to the total runtime.
By contrast, the quantum-inspired Ising-machine simulations proposed in this work, achieve short wall-clock times per energy evaluation, 131-3 s on GPUs. In particular, total potential energy computation for H2 required only $1.2$ s per bond-length parameter $r$ (Fig. 4), while for H2O the simulator required $2.4$ s per parameter $r$ (Fig. 5). Crucially, this highlights a distinct practical advantage of the GPU-based Ising approach by bypassing the queueing latencies and programming overheads inherent to current cloud-based QPUs; we achieve superior computational throughput and immediate scalability.
@@
H2 VQE 79 IBMQ Manila $334\pm 40$ s End-to-end TTS includes queue, compilation, execution, and readout.
H2 VQE 79 AQT Marmot $13{,}297\pm 625$ s End-to-end TTS includes queue, compilation, execution, and readout.
H2O VQE 91 IBMQ Kolkata $5,400$ s Quantum Computation Time.
H2O XBK 78 D-Wave more than 8 s (est.) Estimated from Fig. 3 in the paper (time vs. pre-quadratization qubits).
H2O QCC 78 D-Wave more than 6 s (est.) Estimated from Fig. 3.
@@
H2O Quantum-Inspired Ising Machine GPU 2.4 s Proposed efficient simulation using the Ising Machine approach.
@@ 4 Conclusion
Our numerical analysis, consistent with recent studies that compare the performance of QAIA 82, indicates that CFC variant exhibits superior efficiency compared to other CIM-inspired algorithms 82, 83.
@@
It is noteworthy that our proposed GPU-based algorithm for quantum chemistry calculations can efficiently generate large numbers of samples, enabling the treatment of large-scale scientific problems with minimal time and resource overhead due to its inherent parallelism 49, 48, 52, 53. This high-throughput sampling not only facilitates exploration of Ising-form problems but also provides a powerful framework for investigating electron configurations in molecular systems to identify their stable states. With this physics-inspired, GPU-accelerated approach, the exploration of electronic structures in complex molecules can be performed more efficiently, offering promising applications in material design—such as optimizing catalytic surfaces, studying superconducting materials, and engineering novel semiconductors—and in drug discovery, for instance, in predicting binding affinities, screening molecular conformations, and modeling reaction pathways in biomolecular systems.
@@
Most importantly, in contrast to purely digital optimization approaches, the physical realization of CIMs offers an analog computational paradigm with distinct advantages for quantum chemistry applications. CIMs evolve continuous dynamical variables, such as optical field amplitudes, thereby avoiding the numerical precision limitations and significant memory overhead associated with high-accuracy digital computation. Furthermore, intrinsic quantum noise in optical parametric oscillator networks provides a physically meaningful source of stochasticity, enabling efficient exploration of complex energy landscapes and reducing the likelihood of stagnation in local minima 84. Although the present study employed SimCIM, it is noteworthy that hardware-realized CIMs are expected to offer additional advantages, including faster physical convergence, enhanced scalability, and improved sampling fidelity due to genuine quantum fluctuations. When combined with quantum-inspired algorithms such as CFC, SFC, CAC, and dSB variants as demonstrated in this work, CIM-based hybrid solvers present a scalable and physics-informed alternative to both classical heuristics and hardware-restricted quantum annealers. Their demonstrated ability to approximate molecular ground-state energies using modest computational resources highlights their potential as a key component of future analog–digital hybrid computing frameworks in quantum chemistry.
@@
During the preparation of this work, Takesue et al. 85 demonstrated that a CIM based on degenerate optical parametric oscillators can efficiently solve classically hard NP problems, such as finding maximum independent sets in large graphs, by mapping them onto an Ising Hamiltonian and exploiting the machine’s intrinsic parallelism and energy minimization dynamics. This success story provides a testimony that photonic-based, quantum-inspired systems can efficiently explore complex energy landscapes, a capability directly relevant to quantum chemistry, where molecular Hamiltonians can be encoded into Ising-type models to determine ground-state energies. Leveraging CIMs thus enables systematic exploration of electronic configurations, outperforming traditional digital algorithms in variational molecular simulations.
@@
Building on this foundation, future research can enhance convergence and accuracy through optimized parameter tuning, noise mitigation, and adaptive sampling, while integration with machine learning may enable automated optimization. Embedding these methods in hybrid quantum–classical workflows could further improve efficiency. Extending the approach to larger, chemically relevant systems, including complex biomolecules, remains an important opportunity, and the hybrid algorithm could also be adapted to compute excited-state energy profiles 86, broadening its applicability in quantum chemistry, materials science, and drug discovery.